Noventis Corporation prepared the following estimates for the four quarters of the current year:

First
Quarter
Second
Quarter
Third
Quarter
Fourth
Quarter
Sales $ 1,300,000 $ 1,560,000 $ 1,820,000 $ 2,080,000
Cost of goods sold 430,000 510,000 580,000 630,000
Administrative costs 310,000 185,000 190,000 200,000
Advertising costs 0 100,000 0 0
Executive bonuses 0 0 0 76,000
Provision for bad debts 0 0 0 64,000
Annual maintenance costs 72,000 0 0 0

Additional Information

  • First-quarter administrative costs include the $160,000 annual insurance premium.
  • Advertising costs paid in the second quarter relate to television advertisements that will be broadcast throughout the entire year.
  • No special items affect income during the year.
  • Noventis estimates an effective income tax rate for the year of 40 percent.
  1. Assuming that actual results do not vary from the estimates provided, determine the amount of net income to be reported each quarter of the current year.

  2. Assume that actual results do not vary from the estimates provided except for that in the third quarter, the estimated annual effective income tax rate is revised downward to 38 percent. Determine the amount of net income to be reported each quarter of the current year
Answer #1

Net income to be reported Amount in $

Particulars Qtr 1 Qtr 2 Qtr 3 Qtr 4
Sales 1300,000 1560,000 1820,000 2080,000
Less: Cost of goods sold (430,000) (510,000) (580,000) (630,000)
Gross Profit 870,000 1050,000 1240,000 1450,000
Less: Administrative Costs (Excl annual insurance premium) (150,000) (185,000) (190,000) (200,000)
Less: Annual insurance premium (160,000/4) (40,000) (40,000) (40,000) (40,000)
Less:Advertising costs (To be apportioned over 4 quarters - 100,000 / 4) (25,000) (25,000) (25,000) (25,000)
Less: Executive bonuses (76,000)
Less: Provision for bad debts (64,000)
Less: Annual maintenance costs (To be apportioned over the year - 72,000/4) (18,000) (18,000) (18,000) (18,000)
Profit before Tax 637,000 782,000 967,000 278,000
Less: Income tax expense (254,800) (312,800) (339,080) (105,640)
Net Income after Tax 382,200 469,200 627,920 172,360

Note: Tax for 3 rd quarter

Cumulative Income upto 3rd quarter = 637,000 + 782,000 + 967,000 = $ 23,86,000

Tax expense @ 38% = (23,86,000*38%) = $ 906,680

Less: Cumulative tax expense upto 2nd quarter = $ 254800+$312800= ($567,600)

Balance tax to be recognized in 3rd quarter = $ 339,080